OpenStudy (anonymous):

Find dy/dx by implicit differentation: 5+3x=sin(xy^7) I keep getting... (3-y^7cos(xy^7))/(7xycos(xy^7)) and it's wrong can someone help me?

OpenStudy (anonymous):

differentiate 5+3x giving 3 now sin(xy^7) gives cos(xy^7) * (y^7 + (dy/dx)x7y^6) so 3 = cos(xy^7) * (y^7 + (dy/dx)7xy^6) dy/dx = ([3/ cos(xy^7)] - y^7) / [7xy^6]

OpenStudy (anonymous):

Oh okay, thanks for showing the steps so I could see what I got wrong...

OpenStudy (anonymous):

5+3x=sin(xy7) taking derivatives on both sides 3=cos(xy7)(x7y6.dy/dx+y7) 3-y7cos(xy7)=cos(xy7)(7xy6.dy/dx) dy/dx=3-y7cos(xy7)/cos(xy7)7xy6 that the answer i think so.....the value with y is a power of "y"

OpenStudy (anonymous):

Monroe......my answer is more reliable

OpenStudy (anonymous):

I got it, thanks for the help.

OpenStudy (anonymous):

ok...welcome

OpenStudy (anonymous):

lol its the same thing that i did

OpenStudy (anonymous):

what makes it more reliable ?!

OpenStudy (anonymous):

no...your answer is a bit different.....

OpenStudy (anonymous):

no its not .. you just multiplied numerator and denominator by cos(xy^7)

OpenStudy (anonymous):

look at my brackets.

OpenStudy (anonymous):

oh yes now i see it.......sorry for that....and welldone